# Human-machine social systems

### v2

@@ Abstract
From fake social media accounts and generative-AI chatbots to financial trading algorithms and self-driving vehicles, robots, bots, and algorithms are proliferating and permeating our communication channels, social interactions, economic transactions, and transportation arteries. Networks of multiple interdependent and interacting humans and autonomous machines constitute complex social systems where the collective outcomes cannot be deduced from either human or machine behavior alone. Under this paradigm, we review recent research from across a range of disciplines and identify general dynamics and patterns in situations of competition, coordination, cooperation, contagion, and collective decision-making, with context-rich examples from high-frequency trading markets, a social media platform, an open-collaboration community, and a discussion forum. To ensure more robust and resilient human-machine communities, researchers should study them using complex-system methods, engineers should explicitly design AI for human-machine and machine-machine interactions, and regulators should govern the ecological diversity and social co-evolution of humans and machines.
Introduction
@@
Robotic trains and cars drive us around, auction bots outbid us for purchases, ChatGPT answers our questions, while social media bots feed us with dubious facts and news. Modern society is a complex human-machine social system in which machines are becoming more numerous, human interactions with machines – more frequent, and machine-machine interactions – more consequential. With recent advances in generative AI models, the existential threat of unexplainable and uncontrollable general AI is looming large again 1, 2, 3. However, when they are numerous and interdependent, even simple unintelligent artificial agents can produce unintended and potentially undesirable outcomes. If we want to prevent financial crashes, improve road safety, preserve market competition, increase auction market efficiency, and reduce misinformation, it is no longer sufficient to understand humans – we need to consider machines, understand how humans and machines interact, and how the collective behavior of systems of humans and machines can be predicted. We require a “new sociology of humans and machines.”
@@
This narrative review synthesizes research and ideas related to social systems composed of multiple autonomous yet interacting and interdependent humans and machines such as algorithms, bots, and robots (Fig. 1). Similar to the conceptualizations of socio-technical systems 4, actor-network theory 5, 6, cyber-physical social systems 7, 8, social machines 9, 10, 11, and human-machine networks 12, 13, our principal assumption is that humans and machines form a single social system. In contrast, we do not approach machines as a single medium or entity – “technology” – but emphasize their multiplicity, independence, and heterogeneity and their interactions 14. Thus, we target machines capable of social behavior, namely, a behavior that influences and is influenced by the behavior of other actors (human or machine) in the system. Further, we approach the aggregates as complex social systems where network effects and nonlinear dynamic processes produce collective outcomes that cannot be necessarily deduced from individual preferences and behavior alone15. Our conceptualization extends and complements the ecological approach to studying machine behavior 16, the “hybrid collective intelligence perspective” 17, and the budding field of Social AI 18. We aim to offer a conceptual overview of the topic. Hence, we do not utilize a systematic literature search strategy and instead present selected examples to showcase the new conceptualization 19, 20, 21.
Figure 1: Human-machine social systems include multiple autonomous algorithms, bots, or robots that interact among themselves and with humans in groups and networks. Existing fields tend to either approach machines as media or interfaces, not autonomous actors or agents, or focus on their cognition and decision-making, not group interactions with humans. We call for a new sociology of humans and machines. Researchers should study the human behavior, machine behavior, and the human-human, human-machine, and machine-machine interactions simultaneously in these complex systems, AI engineers should design explicitly for human-machine and machine-machine interactions, and policymakers and regulators should plan for and regulate their ecological diversity and social co-evolution.
@@
We use the term “machines” to refer to various computational artifacts that exist in a multidimensional space ranging on embodiment (from physical devices such as humanoid robots to bots and algorithms that exist in digital space only), algorithmic sophistication (from simple expert systems that use pre-defined if-else rules to generative deep-learning models that learn from data in real-time), autonomy (from fully managed to fully autonomous), sociality (e.g., from infrastructure to human peers), multiplicity (from single large models like GPT4 to many simple models like those used in ensemble learning), heterogeneity (from the coordinated behaviors of bot farms to the diverse strategies of competing trading algorithms), and generality (from highly specialized chess bots to open-ended problem-solvers like LLMs). Here, we restrict our attention to multiple, autonomous, and social artificial agents that operate and interact in the same social environment as humans. Although potentially relevant, in this work, we do not discuss human-assisted bots and bot-assisted humans known as “cyborgs” 22, 23, coordinated bot accounts known as botnets, swarm-bots, and bot farms 24, 25, 26, 27, algorithms that perform background infrastructure-related tasks such as recommender systems 28, 29, crawlers, indexers, and scrapers 30, 31, smart electricity grids 32, and traffic light control systems 33, as well as hybrid traffic flows 34, 35. These systems have been extensively discussed elsewhere. We acknowledge that we may be excluding other interesting cases, including presently out-of-scope cases that future technological advancements may ultimately align with our taxonomy of “autonomous” and “social” machines.
@@
Like humans, the machines we consider exhibit diverse goal-oriented behavior shaped by information and subject to constraints. However, the actual cognition and behavior of machines differ from those of humans. Machines’ behavior tends to be predictable and persistent 36, with higher precision and faster execution 37, better informed with access to global information 38, and less adaptable and susceptible to influence 39, 40, 41. In contrast, humans tend to be limited to local information, satisfice, act with errors, learn and adapt, succumb to social influence and peer pressure, yet also exhibit opinion stubbornness and behavioral inertia; on occasions, they may also use metacognition and revise their own perceptual and decision-making models. Humans often exhibit cognitive biases due to limited information processing capacity, bounded rationality, reliance on heuristics, vestiges of evolutionary adaptation, and emotional motivations 42, 43, and algorithms trained on data generated by humans may reproduce these biases 44, 45. Research on human-like general AI aims to erase the cognitive and behavioral differences between humans and machines, while work on human-competitive AI strives for superintelligence that is faster, smarter, and more precise than humans’ 46, 47. Either way, humans will remain distinct from machines in the near future.
@@
Research from the CASA (computers as social actors) paradigm in psychology emphasizes that humans treat and respond to machines similarly to other humans: people reciprocate kind acts by computers 48, treat them as politely as they treat humans 49, consider them as competent, but also apply gender and racial stereotypes to them 50, 51; people also humanize and empathize with machines, experiencing distress when witnessing the mistreatment of a robot 52, 53.
@@
Nevertheless, there are visible neurophysiological differences in the brain when humans interact with robots 54, 55, likely because humans do not attribute agency and morals to them 56, 57. AI is perceived to have lower intentional capacity, lack self-interest, and be more unbiased than humans. Consequently, humans exhibit a narrower emotional spectrum with machines than with humans, reacting with lower and flatter levels of social emotions such as gratitude, anger, pride, and a sense of fairness 58, 59, 60, 61, yet judging machines more harshly when they commit mistakes, cause harm, or incur losses 62, 63. Further, humans behave more rationally and selfishly with machines, cooperating and sharing less and demanding and exploiting more 64, 65, 66, 67. People would design a machine to be more cooperative than they are themselves 68 but act pro-socially towards it only if it is more human-like 69, or if it benefits another human 70. Compared to a single person, small groups of people are even more likely to exhibit competitive behavior and bullying toward robots 71. Despite this intergroup bias, humans are still susceptible to machine influence when making decisions or solving problems 72. Robots can cause both informational and normative conformity in people 73, 74 and AI and ChatGPT can corrupt humans’ moral judgment and make them follow unethical advice 75, 76. Humans tend to trust algorithmic advice more than advice coming from a another human or a human crowd 77, 78 but may also avoid it if they perceive a threat to their decision control or a lack of understanding and cognitive compatibility 79, 80.
Figure 2: Collective outcomes in human-machine social systems differ from those in human-only systems. First-order differences stem from the fact that machines behave differently from humans; thus, in social systems with covert artificial agents, even if humans, unaware of the presence of machines, do not change their behavior, the collective outcomes will differ simply because machines act differently. Second-order differences occur because humans interact differently with machines than they do with other humans and because machines interact not only with humans but also with each other. Third-order differences are due to the interdependence and mutual influence between the two types of actors and their interactions: simply suspicion or awareness of machine presence can change human behavior while interacting with a machine and observing machine-machine interactions can influence how humans act toward each other.
Table 1: Types, examples, and collective outcomes of human-machine social systems. Boxes 1-4 present more context for four of the examples: high-frequency trading markets, Twitter, Wikipedia, and Reddit. These four communities are clearly defined, relatively large, and well-studied and exemplify situations of market competition, contagion in political communication, cooperation and coordination, and collective action, respectively.
Situation Models General examples H-M examples H-M collective outcomes
Competition Zero-sum game Auction markets Buyer-seller markets Oligopoly market games Competitions Contests Auctions Markets High-frequency trading markets Pricing algorithms Online auctions with sniping algorithms Cheating bots in multiplayer games + Increase efficiency by improving liquidity and price discovery – Increase volatility by causing price spikes and crashes – Increase in consumer prices from algorithmic collusion – Decrease human activity
Coordination Coordination game Stag Hunt game Battle of the Sexes Chicken game Graph Coloring game Conventions Technological standards Communication technology Transportation Supply chain management Traffic with autonomous vehicles Wikipedia editors Humanoid robots in warehouses and manufacturing + Improve coordination by introducing random behavior – Worsen coordination by failing to adapt
Cooperation Prisoner’s Dilemma Public Goods game Dictator game Ultimatum game Trust game Team collaboration Mutual aid Caring robots Personal AI assistants Chatbots Wikipedia editors Reddit moderators Non-player characters in multiplayer games + Increase cooperation + Increase efficiency by handling large task volumes + Increased forecasting accuracy – Decrease efficiency by introducing new types of workload
Contagion Epidemiological models Threshold models of contagion Communicative disease Information Innovations Twitter Reddit – Increase spread of misinformation, opinion polarization, verbal conflict + Increase human activity and engagement
Collective decision making Vote aggregation Active learning Crowdsourcing Prediction markets Voting systems Cultural evolution Clinical diagnosis Citizen science Content moderation Hybrid forecasting + Increase innovation and accuracy by introducing diversity – Decrease human activity and engagement
@@
In complex social systems, individuals’ behavior and interactions affect the collective outcomes, but the relationship can be fundamentally different from a simple sum or average 81, 82, 83, 84. The collective outcomes in human-machine social systems differ from those in human-only systems because machines behave differently from humans, human-machine (H-M) and machine-machine (M-M) interactions differ from human-human (H-H) interactions, but also the humans, the machines, and their interactions influence each other indirectly (Fig. 2). We synthesize common dynamics and patterns in groups and networks of humans and machines for five different social interaction situations (Table 1).
@@
With more advanced data processing, learning, and optimization capabilities than humans, algorithms are better able to discover arbitrage opportunities and, hence, eliminate mispricing and increase liquidity in markets. Experimental studies show that algorithmic traders can increase market efficiency 85, but possibly at the expense of human traders’ performance 86, 87. Furthermore, algorithmic traders affect human behavior indirectly: with their presence, they make human traders act more rationally, and thus, reduce strategic uncertainty and confusion in the market 88, reducing price bubbles and bringing prices closer to the fundamental value 89.
@@
While perfectly optimizing arbitrage algorithms eliminate mispricings, neither zero-intelligence algorithms that submit random bids without profit maximization 90 nor profit-maximizing agents that update their beliefs from trading history 91 can improve market quality. Meanwhile, manipulator and spoofing algorithms that act to mislead and influence other traders worsen market efficiency 92. Algorithms can also reduce the rationality of professional traders and alienate and drive away amateur ones. In an online cryptocurrency marketplace, traders herd after a bot buys, producing larger buying volumes 93, while in a Chinese peer-to-peer lending platform, automated investment piques inefficient investor scrutiny, increasing bidding duration without improving investment return 94. In online auction markets, naive first-time bidders respond negatively to being outbid by sniping algorithms and become less likely to return to another auction 95. Snipers, which place last-moment bids 96, work mainly because they exploit the naivety of amateur online bidders, who tend to increase their bids incrementally. However, human lack of rationality has its benefits because squatting (placing a high early bid) deters new entrants 97. In fact, sniping algorithms yield a negligibly small 97 or non-existent 98 buyer gain, giving them a net negative impact on the marketplace.
@@
In addition to trading and auction markets, pricing algorithms have become widespread in regular product markets 99, 100 because they either provide recommendations to human pricing managers 101, 102 or entirely dictate pricing for some firms 100, 103. While pricing algorithms can help firms scale and respond to changes in demand, they may also generate anti-competition. Q-learning algorithms learn to set anti-competitive prices without communication in simulations 104, 105, 106, 107, and in experiments, those algorithms are often more collusive than humans in small markets 108 and foster collusion when interacting with humans compared to fully human markets 109. Observational studies of gasoline markets 103 and e-commerce 110, 111 support the experimental evidence. Furthermore, algorithms can weaken competition by providing better demand predictions, thereby stabilizing cartels 112, 113, 114 or by asymmetries in pricing technologies and commitment 115, 116.
@@
The general intuition is that markets with more actors should be more efficient. Thus, one expects enhanced performance from markets populated by algorithms. However, in reality, the beneficial effects of machines are often in balance, crucially depending on the machines’ prevalence, decision speed, and information quality 117, as well as the humans’ experience and expectations.
@@
The problem of coordination requires adopting a strategy identical to or, in some cases, dissimilar from other people’s strategies, as when deciding whether to join a protest, agreeing on a convention such as driving on the right-hand side of the road, adopting a communication technology, or avoiding a crowd or traffic congestion 118, 119.
In human-machine systems, bots could be used to introduce more randomness and movement to steer human groups toward better solutions. Thus, bots acting with small levels of random noise and placed in central locations in a scale-free network decrease the time to coordination, especially when the solutions are hard to find 120. The bots reduce unresolvable conflicts not only in their direct interactions but also in indirect H-H interactions, even when the participants are aware that they are interacting with machines. Bots that are trained on human behavior, however, process information less efficiently and adapt slower, causing hybrid groups playing a cooperative group-formation game to perform worse than human-only and bot-only groups 41.
@@
The problem of cooperation pertains to social dilemma situations where a decision is collectively beneficial but individually costly and risky. Although the economically rational decision in non-repeated anonymous interactions is to free-ride and exploit others’ contributions, people’s actual behavior tends to be informed by norms of reciprocity, fairness, and honesty signaling. Thus, as a result of millennia of evolutionary adaptation, people generally cooperate with each other. If people know they are interacting with bots, however, they cooperate less 65, 66. Yet, since humans reciprocate to and imitate cooperative neighbors, introducing covert, persistently cooperating bots could increase cooperation.
Computer simulations show that persistent prosocial bots favor the emergence of fairness and cooperation 121, 122, with stronger effects when humans are more prone to imitation and bots occupy more central positions in networks with highly heterogeneous connectivity 123.
Just a few under-cover cooperative bots can increase cooperation, especially if the bots are widespread in the network, interacting with many human players rather than concentrated with overlapping sets of partners 124. The reason is that humans wait for someone else to cooperate before they do, but once they observe many cooperators, they become more likely to exploit.
Yet, cooperative bots may sometimes fail to improve cooperation. For instance, hybrid groups with identifiable bots do not perform better than human-only groups 125. When participants are aware of the presence of artificial agents but not their identity, there is a small increase in cooperation of the bots’ direct neighbors but no significant boost in the overall network 126. Similarly, multiple well-dispersed covert bots, whether all-cooperating or reciprocating, fail to improve cooperation 127, although a single overt network engineer bot who suggests connecting cooperators and excluding defectors can successfully do so.
@@
Contagion concerns the spread of information and behaviors, such as memes, slang, fashion, emotions, and opinions in communication networks 128, 129, 130. In contrast to the strategic interdependence under the competition, coordination, and cooperation scenarios, the main mechanism here is social influence: the tendency to rely on information from others to handle uncertainty and to conform to the expectations of others to fit in society 131, 132, 133. In human-machine systems, bots can be remarkably influential at the collective level despite exerting limited direct influence on individuals because, in networks, small effects can produce chain reactions and trigger cascades 134, 135, 136, 137.
This is how social media bots influence public opinion. In agent-based models of belief formation, weak bots do not alienate their followers and their followers’ friends and thus have their message spread farther than messages by more pushy and assertive users 138. In other words, network amplification occurs through bots’ indirect influence precisely because their direct influence on humans is weak, slow, and unobtrusive. If social media bots influence not people’s opinion but their confidence to express it, they can amplify marginal voices by triggering the spiral of silence amongst disagreeing humans 40. The bots are more influential when they are more numerous and connected to central actors. Strategically placed zealot bots can in fact bias voting outcomes and win elections 139.
Bots can also trigger emotional contagion in groups, even though they evoke flatter emotional reactions from individual humans. Humanoid robots can encourage and increase social interactions among older adults within care facilities, between different generations, and for children with ASD 71. In small-team collaborative experiments, a robot’s verbal expressions of vulnerability can show “ripple effects” and make the humans more likely to admit mistakes, console team members, and laugh together 140, engage in social conversations and appreciate the group 141. The reported positive contagion effects, however, were detected when comparing one machine to another 141, 142. Overall, bots are more effective than no bots to influence opinions, behavior, and emotions, but not necessarily more effective than humans. Yet, even when bots have a weak direct influence on humans’ opinions, they can exert significant collective influence via persistence, strategic placement, and sheer numbers.
@@
Collective decision-making involves groups making choices or solving problems by combining individual opinions. It impacts social phenomena as diverse as team collaboration, voting, scientific innovation, and cultural evolution 143. Originating with Galton’s work on estimation tasks 144, the “wisdom of crowds” concept suggests that a crowd’s aggregated estimate is often more accurate than any individual’s, or sometimes even experts’ 145. Crowds perform better when individual opinions are either independent or diverse 146, while social interaction can hinder 147, 148, 149 or improve 150, 151 collective performance. In human-machine systems, algorithms introduce diversity and can thus improve decision-making.
@@
An analysis of professional Go players’ moves over 71 years suggests that AlphaGo, the AI program Google DeepMind introduced in 2016, led human players to novel strategies and improved their decision-making 152, 153, 154. AlphaGo’s decisions, untethered by human bias, sparked human innovation in this game. However, the positive influence of machine-human social learning on problem-solving may be limited. When algorithms are introduced in chains of humans engaged in sequential problem solving, the innovative solutions benefit immediate followers, but team accuracy does not have lasting effects because humans are more likely to replicate human solutions than algorithmic ones 155. Similarly, in a team prediction task, an algorithm maintaining group diversity by promoting minority opinions improves individual accuracy, but the effects dissipate for team accuracy 156.
@@
The area of hybrid intelligence investigates how and when to combine human and algorithmic decision-making 157, 17 and includes research on active and interactive learning and human-in-the-loop algorithms 158, with applications in clinical decision-making, where combining clinician and algorithmic judgments can improve cancer diagnoses 159, 160, and citizen science, where combining crowd-based with machine classifications can improve accuracy. On Zooniverse, a prominent citizen science platform, this hybrid approach found supernovae candidates among Pan-STARRS images more effectively than humans or machines alone 161 but damaged citizen scientists’ retention 162, 163, suggesting a trade-off between efficiency and volunteer engagement. Ultimately, the deployment of machines could further marginalize certain groups of volunteers 164, and with fewer volunteers, AI’s performance could diminish.

### v3

# A new sociology of humans and machines

@@ Abstract
From fake social media accounts and generative-AI chatbots to trading algorithms and self-driving vehicles, robots, bots, and algorithms are proliferating and permeating our communication channels, social interactions, economic transactions, and transportation arteries. Networks of multiple interdependent and interacting humans and intelligent machines constitute complex social systems where the collective outcomes cannot be deduced from either human or machine behavior alone. Under this paradigm, we review recent research and identify general dynamics and patterns in situations of competition, coordination, cooperation, contagion, and collective decision-making, with context-rich examples from high-frequency trading markets, a social media platform, an open-collaboration community, and a discussion forum. To ensure more robust and resilient human-machine communities, we require a new sociology of humans and machines. Researchers should study these communities using complex-system methods, engineers should explicitly design AI for human-machine and machine-machine interactions, and regulators should govern the ecological diversity and social co-development of humans and machines.
@@
Robotic trains and cars drive us around, auction bots outbid us for purchases, ChatGPT answers our questions, while social media bots feed us with dubious facts and news. Modern society is a complex human-machine social system in which machines are becoming more numerous, human interactions with machines – more frequent, and machine-machine interactions – more consequential. With recent advances in generative AI models, the existential threat of unexplainable and uncontrollable general AI is looming large again 1, 2, 3. However, when they are numerous and interdependent, even simple unintelligent artificial agents can produce unintended and potentially undesirable outcomes. If we want to prevent financial crashes, improve road safety, preserve market competition, increase auction market efficiency, and reduce misinformation, it is no longer sufficient to understand humans – we need to consider machines, understand how humans and machines interact, and how the collective behavior of systems of humans and machines can be predicted. We require a new sociology of humans and machines.
@@
This Perspective synthesizes research and ideas related to social systems composed of multiple autonomous yet interacting and interdependent humans and machines such as algorithms, bots, and robots (Fig. 1). Similar to the conceptualizations of socio-technical systems 4, actor-network theory 5, 6, cyber-physical social systems 7, 8, social machines 9, 10, 11, and human-machine networks 12, 13, our principal assumption is that humans and machines form a single social system. In contrast, we do not approach machines as a single medium or entity – “technology” – but emphasize their multiplicity, independence, and heterogeneity and their interactions 14. Further, we approach the aggregates as complex social systems where network effects and nonlinear dynamic processes produce collective outcomes that cannot be necessarily deduced from individual preferences and behavior alone15. Our conceptualization extends and complements the ecological approach to studying machine behavior 16, the “hybrid collective intelligence perspective” 17, and the budding field of Social AI 18. We aim to offer a conceptual overview of the topic. Hence, we do not utilize a systematic literature search strategy and instead present selected examples to showcase the new conceptualization 19, 20, 21.
@@
We use the term “machines” to refer to a range of computational artifacts: embodied in physical devices such as humanoid robots or existing only in digital space, such as bots and algorithms, and varying in sophistication from simple expert systems that use pre-defined if-else rules to generative deep-learning models that learn from data in real-time. Like humans, the machines we consider exhibit diverse goal-oriented behavior shaped by information and subject to constraints. However, the actual cognition and behavior of machines differ from those of humans. Machines’ behavior tends to be predictable and persistent 22, with higher precision and faster execution 23, better informed with access to global information 24, and less adaptable and susceptible to influence 25, 26, 27. In contrast, humans tend to be limited to local information, satisfice, act with errors, learn and adapt, succumb to social influence and peer pressure, yet also exhibit opinion stubbornness and behavioral inertia; on occasions, they may also use metacognition and revise their own perceptual and decision-making models. Humans often exhibit cognitive biases due to limited information processing capacity, bounded rationality, reliance on heuristics, vestiges of evolutionary adaptation, and emotional motivations 28, 29, and algorithms trained on data generated by humans may reproduce these biases 30, 31. Research on human-like general AI aims to erase the cognitive and behavioral differences between humans and machines, while work on human-competitive AI strives for superintelligence that is faster, smarter, and more precise than humans’ 32, 33. Either way, humans will remain distinct from machines in the near future.
@@
Research from the CASA (computers as social actors) paradigm in psychology emphasizes that humans treat and respond to machines similarly to other humans: people reciprocate kind acts by computers 34, treat them as politely as they treat humans 35, consider them as competent, but also apply gender and racial stereotypes to them 36, 37; people also humanize and empathize with machines, experiencing distress when witnessing the mistreatment of a robot 38, 39.
@@
Nevertheless, there are visible neurophysiological differences in the brain when humans interact with robots 40, 41, likely because humans do not attribute agency and morals to them 42, 43. AI is perceived to have lower intentional capacity, lack self-interest, and be more unbiased than humans. Consequently, humans exhibit a narrower emotional spectrum with machines than with humans, reacting with lower and flatter levels of social emotions such as gratitude, anger, pride, and a sense of fairness 44, 45, 46, 47, yet judging machines more harshly when they commit mistakes, cause harm, or incur losses 48, 49. Further, humans behave more rationally and selfishly with machines, cooperating and sharing less and demanding and exploiting more 50, 51, 52, 53. People would design a machine to be more cooperative than they are themselves 54 but act pro-socially towards it only if it is more human-like 55, or if it benefits another human 56. Compared to a single person, small groups of people are even more likely to exhibit competitive behavior and bullying toward robots 57. Despite this intergroup bias, humans are still susceptible to machine influence when making decisions or solving problems 58. Robots can cause both informational and normative conformity in people 59, 60 and AI and ChatGPT can corrupt humans’ moral judgment and make them follow unethical advice 61, 62. Humans tend to trust algorithmic advice more than advice coming from a another human or a human crowd 63, 64 but may also avoid it if they perceive a threat to their decision control or a lack of understanding and cognitive compatibility 65, 66.
@@
In complex social systems, individuals’ behavior and interactions affect the collective outcomes, but the relationship can be fundamentally different from a simple sum or average 67, 68, 69, 70. The collective outcomes in human-machine social systems differ from those in human-only systems because machines behave differently from humans, human-machine (H-M) and machine-machine (M-M) interactions differ from human-human (H-H) interactions, but also the humans, the machines, and their interactions influence each other indirectly (Fig. 2). We synthesize common dynamics and patterns in groups and networks of humans and machines for five different social interaction situations (Table 1).
@@
With more advanced data processing, learning, and optimization capabilities than humans, algorithms are better able to discover arbitrage opportunities and, hence, eliminate mispricing and increase liquidity in markets. Experimental studies show that algorithmic traders can increase market efficiency 71, but possibly at the expense of human traders’ performance 72, 73. Furthermore, algorithmic traders affect human behavior indirectly: with their presence, they make human traders act more rationally, and thus, reduce strategic uncertainty and confusion in the market 74, reducing price bubbles and bringing prices closer to the fundamental value 75.
@@
While perfectly optimizing arbitrage algorithms eliminate mispricings, neither zero-intelligence algorithms that submit random bids without profit maximization 76 nor profit-maximizing agents that update their beliefs from trading history 77 can improve market quality. Meanwhile, manipulator and spoofing algorithms that act to mislead and influence other traders worsen market efficiency 78. Algorithms can also reduce the rationality of professional traders and alienate and drive away amateur ones. In an online cryptocurrency marketplace, traders herd after a bot buys, producing larger buying volumes 79, while in a Chinese peer-to-peer lending platform, automated investment piques inefficient investor scrutiny, increasing bidding duration without improving investment return 80. In online auction markets, naive first-time bidders respond negatively to being outbid by sniping algorithms and become less likely to return to another auction 81. Snipers, which place last-moment bids 82, work mainly because they exploit the naivety of amateur online bidders, who tend to increase their bids incrementally. However, human lack of rationality has its benefits because squatting (placing a high early bid) deters new entrants 83. In fact, sniping algorithms yield a negligibly small 83 or non-existent 84 buyer gain, giving them a net negative impact on the marketplace.
@@
In addition to trading and auction markets, pricing algorithms have become widespread in regular product markets 85, 86 because they either provide recommendations to human pricing managers 87, 88 or entirely dictate pricing for some firms 86, 89. While pricing algorithms can help firms scale and respond to changes in demand, they may also generate anti-competition. Q-learning algorithms learn to set anti-competitive prices without communication in simulations 90, 91, 92, 93, and in experiments, those algorithms are often more collusive than humans in small markets 94 and foster collusion when interacting with humans compared to fully human markets 95. Observational studies of gasoline markets 89 and e-commerce 96, 97 support the experimental evidence. Furthermore, algorithms can weaken competition by providing better demand predictions, thereby stabilizing cartels 98, 99, 100 or by asymmetries in pricing technologies and commitment 101, 102.
@@
The general intuition is that markets with more actors should be more efficient. Thus, one expects enhanced performance from markets populated by algorithms. However, in reality, the beneficial effects of machines are often in balance, crucially depending on the machines’ prevalence, decision speed, and information quality 103, as well as the humans’ experience and expectations.
@@
The problem of coordination requires adopting a strategy identical to or, in some cases, dissimilar from other people’s strategies, as when deciding whether to join a protest, agreeing on a convention such as driving on the right-hand side of the road, adopting a communication technology, or avoiding a crowd or traffic congestion 104, 105.
In human-machine systems, bots could be used to introduce more randomness and movement to steer human groups toward better solutions. Thus, bots acting with small levels of random noise and placed in central locations in a scale-free network decrease the time to coordination, especially when the solutions are hard to find 106. The bots reduce unresolvable conflicts not only in their direct interactions but also in indirect H-H interactions, even when the participants are aware that they are interacting with machines. Bots that are trained on human behavior, however, process information less efficiently and adapt slower, causing hybrid groups playing a cooperative group-formation game to perform worse than human-only and bot-only groups 27.
@@
The problem of cooperation pertains to social dilemma situations where a decision is collectively beneficial but individually costly and risky. Although the economically rational decision in non-repeated anonymous interactions is to free-ride and exploit others’ contributions, people’s actual behavior tends to be informed by norms of reciprocity, fairness, and honesty signaling. Thus, as a result of millennia of evolutionary adaptation, people generally cooperate with each other. If people know they are interacting with bots, however, they cooperate less 51, 52. Yet, since humans reciprocate to and imitate cooperative neighbors, introducing covert, persistently cooperating bots could increase cooperation.
Computer simulations show that persistent prosocial bots favor the emergence of fairness and cooperation 107, 108, with stronger effects when humans are more prone to imitation and bots occupy more central positions in networks with highly heterogeneous connectivity 109.
Just a few under-cover cooperative bots can increase cooperation, especially if the bots are widespread in the network, interacting with many human players rather than concentrated with overlapping sets of partners 110. The reason is that humans wait for someone else to cooperate before they do, but once they observe many cooperators, they become more likely to exploit.
Yet, cooperative bots may sometimes fail to improve cooperation. For instance, hybrid groups with identifiable bots do not perform better than human-only groups 111. When participants are aware of the presence of artificial agents but not their identity, there is a small increase in cooperation of the bots’ direct neighbors but no significant boost in the overall network 112. Similarly, multiple well-dispersed covert bots, whether all-cooperating or reciprocating, fail to improve cooperation 113, although a single overt network engineer bot who suggests connecting cooperators and excluding defectors can successfully do so.
@@
Contagion concerns the spread of information and behaviors, such as memes, slang, fashion, emotions, and opinions in communication networks 114, 115, 116. In contrast to the strategic interdependence under the competition, coordination, and cooperation scenarios, the main mechanism here is social influence: the tendency to rely on information from others to handle uncertainty and to conform to the expectations of others to fit in society 117, 118, 119. In human-machine systems, bots can be remarkably influential at the collective level despite exerting limited direct influence on individuals because, in networks, small effects can produce chain reactions and trigger cascades 120, 121, 122, 123.
This is how social media bots influence public opinion. In agent-based models of belief formation, weak bots do not alienate their followers and their followers’ friends and thus have their message spread farther than messages by more pushy and assertive users 124. In other words, network amplification occurs through bots’ indirect influence precisely because their direct influence on humans is weak, slow, and unobtrusive. If social media bots influence not people’s opinion but their confidence to express it, they can amplify marginal voices by triggering the spiral of silence amongst disagreeing humans 26. The bots are more influential when they are more numerous and connected to central actors. Strategically placed zealot bots can in fact bias voting outcomes and win elections 125.
Bots can also trigger emotional contagion in groups, even though they evoke flatter emotional reactions from individual humans. Humanoid robots can encourage and increase social interactions among older adults within care facilities, between different generations, and for children with ASD 57. In small-team collaborative experiments, a robot’s verbal expressions of vulnerability can show “ripple effects” and make the humans more likely to admit mistakes, console team members, and laugh together 126, engage in social conversations and appreciate the group 127. The reported positive contagion effects, however, were detected when comparing one machine to another 127, 128. Overall, bots are more effective than no bots to influence opinions, behavior, and emotions, but not necessarily more effective than humans. Yet, even when bots have a weak direct influence on humans’ opinions, they can exert significant collective influence via persistence, strategic placement, and sheer numbers.
@@
Collective decision-making involves groups making choices or solving problems by combining individual opinions. It impacts social phenomena as diverse as team collaboration, voting, scientific innovation, and cultural evolution 129. Originating with Galton’s work on estimation tasks 130, the “wisdom of crowds” concept suggests that a crowd’s aggregated estimate is often more accurate than any individual’s, or sometimes even experts’ 131. Crowds perform better when individual opinions are either independent or diverse 132, while social interaction can hinder 133, 134, 135 or improve 136, 137 collective performance. In human-machine systems, algorithms introduce diversity and can thus improve decision-making.
@@
An analysis of professional Go players’ moves over 71 years suggests that AlphaGo, the AI program Google DeepMind introduced in 2016, led human players to novel strategies and improved their decision-making 138, 139, 140. AlphaGo’s decisions, untethered by human bias, sparked human innovation in this game. However, the positive influence of machine-human social learning on problem-solving may be limited. When algorithms are introduced in chains of humans engaged in sequential problem solving, the innovative solutions benefit immediate followers, but team accuracy does not have lasting effects because humans are more likely to replicate human solutions than algorithmic ones 141. Similarly, in a team prediction task, an algorithm maintaining group diversity by promoting minority opinions improves individual accuracy, but the effects dissipate for team accuracy 142.
@@
The area of hybrid intelligence investigates how and when to combine human and algorithmic decision-making 143, 17 and includes research on active and interactive learning and human-in-the-loop algorithms 144, with applications in clinical decision-making, where combining clinician and algorithmic judgments can improve cancer diagnoses 145, 146, and citizen science, where combining crowd-based with machine classifications can improve accuracy. On Zooniverse, a prominent citizen science platform, this hybrid approach found supernovae candidates among Pan-STARRS images more effectively than humans or machines alone 147 but damaged citizen scientists’ retention 148, 149, suggesting a trade-off between efficiency and volunteer engagement. Ultimately, the deployment of machines could further marginalize certain groups of volunteers 150, and with fewer volunteers, AI’s performance could diminish.